OpenStudy (anonymous):

Solve for x:

OpenStudy (anonymous):

\[\sqrt{\log_{2}x-1}-1/2\log_{2}x^{3}+2>0\] Ans. x belongs to [2,4)

OpenStudy (anonymous):

let y=log2(x) and solve for y

OpenStudy (lgbasallote):

lol wow another long one..hmm

OpenStudy (anonymous):

anyone solving ??

OpenStudy (anonymous):

@nitz

OpenStudy (anonymous):

@lgbasallote Do you know the solution to it??

OpenStudy (anonymous):

is this the inequality: \(\large \sqrt{log_2x-1}-\frac{1}{2}log_2x^3+2>0 \)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\(\large \sqrt{log_2x-1}-\frac{1}{2}log_2x^3+2>0 \) \(\large \sqrt{log_2x-1}-\frac{3}{2}log_2x+2>0 \) \(\large \sqrt{log_2x-1}>\frac{3}{2}log_2x-2 \) \(\large log_2x-1>(\frac{3}{2}log_2x-2)^2 \) looks like a quadratic form from here... can you take it from here by using what @mukushla said?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

that inequality will look like this: \(\large y-1>(\frac{3}{2}y-2)^2 \)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw...:)